# Singularity resolution and inflation from an infinite tower of regularized curvature corrections

### v1

@@ Abstract
Recent work has shown that regular black holes solutions exist [11] and can form dynamically [12, 13] in higher dimensions, $D\geq 5$, within the framework of quasi-topological gravity [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] when accounting for an infinite tower of curvature corrections. This represents a major step forward in tackling one of the key challenges in theoretical physics: identifying a mechanism to resolve spacetime singularities [24]. Unlike many phenomenological regular black hole models, this approach avoids the need for ad hoc exotic matter or theories that impose fine-tuned relationships between a solution’s integration constants and the theory’s parameters, as often seen in non-linear electrodynamics models [25, 26, 27]. In cosmological settings, it has been shown that, in a similar framework, infinite towers of curvature corrections [23] can replace the initial singularity with an inflationary period [28]. This framework is known as “geometric inflation” [28, 29, 30, 31, 32, 33]. Unfortunately, these theories generally lead to field equations of higher-order in derivatives, except in specific backgrounds, making them susceptible to Ostrogradsky instabilities [34, 35], and quasi-topological gravities do not exist in four-dimensions [18].
@@
In recent years, there has been a surge of interest in the quadratic Lovelock invariant – the Gauss-Bonnet term – and the possibility of a non-trivial Gauss-Bonnet–corrected theory of gravity in four dimensions. This idea was first introduced in Ref. [38] through a singular rescaling of the Gauss-Bonnet coupling constant in a four-dimensional limit of the higher-dimensional theory. However, this singular limit was later shown to be ill-defined [39], leading to the development of well-defined regularizations. These regularizations resulted in specific four-dimensional Horndeski scalar-tensor theories [40, 41, 42, 43, 44], which preserve many solutions and properties of the higher-dimensional Einstein-Gauss-Bonnet theory. As a result, they have been extensively studied; see Ref. [45] for a comprehensive review.
@@
To obtain Lovelock-like corrections at all orders curvature in four dimensions, we apply a well-defined dimensional regularization to each Lovelock invariant, resulting in scalar-tensor theories within the Horndeski class [46, 47]. This regularization method was first introduced in Ref. [48] to recover GR-like dynamics in two dimensions by regularizing the Ricci scalar and was later extended in Refs. [40, 41] to formulate a well-defined Gauss-Bonnet theory in four dimensions. This method relies on the use of two conformally related metrics, $\tilde{g}_{\mu\nu}=e^{-2\phi}g_{\mu\nu}$, and the following limit at each Lovelock order
$\mathcal{L}^{(n)}=\lim_{d\to 2n}\frac{\sqrt{-g}\mathcal{R}^{(n)}-\sqrt{-\tilde{g}}\tilde{\mathcal{R}}^{(n)}}{d-2n},$ (2)
where the expressions inside the limit are to be evaluated in $d$-dimensions, and the limit is taken as $d$ approaches the critical dimension $2n$, where the $n^{\rm th}$ Lovelock invariant becomes topological. This computation for $n=1$ in $2D$, and $n=2$ in $4D$ was performed in detail in Ref. [40], and it can be shown that the limit is well-defined for every $n$, resulting in a scalar-tensor theory, with scalar $\phi$, and second-order equations of motion [49]. Although the limit is taken as $d\to 2n$, the resulting scalar-tensor Lagrangian can be evaluated within a four-dimensional action for any $n$. This four-dimensional theory is inspired by the three-dimensional limits of Gauss-Bonnet gravity and their solutions [50, 51], which, despite not being considered in the Gauss-Bonnet critical dimension $D=4$, lead to solutions similar to those in higher-dimensional Gauss-Bonnet gravity. The final result at each $n$ for the limit (2) is that $\mathcal{L}^{(n)}$ is proportional to a four-dimensional Horndeski Lagrangian with functions given by [49]
@@
$G_{3}^{(n)}=-2^{n}n(2n-3)X^{n-1},$
$G_{4}^{(n)}=2^{n-1}nX^{n-1},$
$G_{5}^{(n)}=-\begin{cases}4\log X,\qquad n=2,\\ 2^{n-1}\frac{n(n-1)}{n-2}X^{n-2},\qquad n>2,\end{cases}$
@@
where $\{c_{n}\}$ are dimensionless coupling constants, $\ell$ is a new length scale, and now $\mathcal{L}^{(n)}$ is the Horndeski Lagrangian with functions given by Eq. (3). This theory falls within the shift-symmetric class of Horndeski theories, implying the existence of a conserved current $J^{\mu}$, whose explicit form is provided in Ref. [52]. The vanishing divergence of this current, $\nabla_{\mu}J^{\mu}=0$, is equivalent to the scalar field equation of motion. Importantly, a sufficient condition to solve the scalar field equation of motion, is $J^{\mu}=0$. The Einstein equations for a generic Horndeski Lagrangian are presented in Refs. [53, 54].
When the whole tower of corrections is considered, the theory described by Eq. (4) can in some cases be re-summed into Horndeski theories with a non-local structure. Some examples are presented in the Supplemental Material.
Figure 1: (Left) Time evolution of $a$ and $H^{2}$ (in units $\ell=1$) for the Friedmann equation in Eq. (9), and in GR, assuming a radiation dominated universe with energy density $\rho=3/(32\pi)$ and scale factor $a=1$ at time $t=1$. In GR, a Big Bang singularity occurs at $t=0$, whereas when an infinite tower of corrections is considered, the singularity is replaced by an inflationary phase. (Right) Slow-roll parameter $-\dot{H}/H^{2}$ as a function of the number of e-folds $N$ (defined such that $N=0$ at the end of inflation), for the theories with couplings given by $c_{n}=(1-(-1)^{n})/(2n)$ (solid line), and $c_{n}=1/n!$ (dashed line).
@@
For the theory (4) on an FLRW background, the only non-trivial component of the shift-symmetry current is $J^{t}$, that at each order $n$, vanishes if the scalar field obeys
@@
where $a_{0}$ is an arbitrary integration constant. Therefore, this profile solves the scalar field equation of motion, regardless of $n$. Using this scalar field profile in the Einstein equations, we obtain the generalized Friedmann equation
@@
We observe that just as in Lovelock gravity [56], the $n^{\text{th}}$ order contribution to the Friedmann equation is proportional to $H^{2n}$. As an example, for the theory with $c_{n}=\left(1-(-1)^{n}\right)/(2n)$, we obtain $F(H^{2})=\tanh^{-1}\left(\ell^{2}H^{2}\right)/\ell^{2}$ such that the Friedmann equation can be written as
@@
In this scenario, as $\rho\to\infty$ in the very early universe [^2] From the continuity equation, we obtain $\rho=\rho_{0}a^{-3(1+\omega)}$, where $\omega$ is the equation of state parameter, $p=\omega\rho$. In the early universe, radiation ($\omega=1/3$) dominates., the spacetime must asymptotically approach a non-singular de Sitter geometry with $H^{2}\to 1/\ell^{2}$. Consequently, the Big Bang singularity is replaced by an inflationary phase in the early universe. This can be observed in Fig. 1 (left) we compare the time evolution of $H^{2}$ and $a$ for a radiation dominated universe in GR, and in the theory with Friedmann equation given in Eq. (9). For late times, the evolution of the universe is well-described by GR. Other couplings we have explored, such as $c_{n}=1$, $c_{n}=1/n$, $c_{n}=1/n!$, and $c_{n}=1/(n-1)!$, also replace the initial singularity with an inflationary phase. In Fig. 1 (right), we plot the slow-roll parameter $-\dot{H}/H^{2}$ as a function of the number of e-folds $N$, for the theories with $F(H^{2})=\tanh^{-1}\left(\ell^{2}H^{2}\right)/\ell^{2}$ and $F(H^{2})=(\exp\left(\ell^{2}H^{2}\right)-1)/\ell^{2}$. We observe an inflationary period ($-\dot{H}/H^{2}<1$) followed by a graceful transition into a radiation dominated universe ($-\dot{H}/H^{2}=2$). Our findings align with those of Refs. [28, 58], who investigated Friedmann equations of the form (7), and generically identified the presence of inflationary phases in the early universe with a natural transition mechanism into standard GR evolution.
Regular black holes. Black holes are predicted to form through the gravitational collapse of massive objects. The simplest scenario for such collapse, initially explored by Oppenheimer and Snyder [59], involves the collapse of a homogeneous and isotropic sphere of pressureless dust. The interior metric describing this collapsing dust, which must be matched to an exterior metric, can in some cases be represented by the FLRW line-element (5) and its dynamics governed by the Friedmann equations. In GR, the dust collapses in finite time to form a singularity and a Schwarzschild black hole. However, when the full tower of corrections in Eq. (4) is taken into account, the Friedmann equations are modified and given by Eq. (7). For the examples studied in the previous section, such as Eq. (9), the collapsing dust never forms a singularity in finite time. This raises questions about whether singularities can form at all through gravitational collapse in this framework.
@@
where the function $F(x)$ was defined in Eq. (8) and $f_{\rm GR}(r)$ in Eq. (11). This leads us to a remarkable conclusion: the class of theories examined in this letter satisfy a Birkhoff-type theorem [^4] To our knowledge, this is the first example of a Birkhoff-type theorem holding for a class of higher-derivative theories beyond GR in four dimensions. However, see Refs. [76, 83, 84, 13, 12] for specific cases in $2+1$ dimensions and higher-dimensional scenarios where a Birkhoff theorem has been established.. Specifically, for the scalar field profile (12), any solution of the form (10) within the theory (4) must be both static and unique, characterized by $N(t,r)=1$ and determined by the solution to Eq. (13). This result can be used to study the collapse of matter and the dynamical formation of black holes, following e.g. the approach of Refs. [13, 12].
As discussed in Refs. [66, 67, 68, 69], for static spacetimes of the form (10), a necessary and sufficient condition for the regularity of all independent components of the Riemann tensor, $R^{\mu\nu}_{\phantom{\mu}\phantom{\nu}\alpha\beta}$, and consequently all curvature invariants constructed from it, is that the Kretschmann scalar remains finite everywhere. By analyzing the Kretschmann scalar, one can show that the singularity at $r=0$ is resolved if the metric function behaves as $f(r)=-r^{2}/\lambda^{2}+\mathcal{O}(r^{3})$ near the origin, where $\lambda$ is a constant with dimensions of length. From Eq. (13), this suggests that resolving the singularity may be possible if the function $F(x)$ develops a pole as $x\to 1/\lambda^{2}$. Since there are infinitely many possible choices of the coefficients $\{c_{n}\}$ that could induce such a pole in $F(x)$, we adopt as a representative example the same couplings used in the previous section, $c_{n}=\left(1-(-1)^{n}\right)/(2n)$, which leads to a pole at $x\to 1/\ell^{2}$. In this case, the solution to Eq. (13) is
@@
where the GR solution is recovered in the limit $\ell\to 0$. The location of the event horizon remains unchanged from the GR case, $r_{H}=(2L^{2}M)^{1/3}$, and the solution is asymptotically AdS. This metric function is everywhere smooth and non-divergent, even when analytically extended to negative values of the coordinate $r$ [70, 71]. The solution is also free of standard curvature singularities as the Kretschmann scalar is everywhere bounded, obeying $\lim_{r\to 0}R_{\mu\alpha\nu\beta}R^{\mu\alpha\nu\beta}=24/\ell^{4}$. Additional solutions which are non-singular at $r=0$ can be found, for instance, in the theories where $c_{n}=1$ or $c_{n}=1/n$.
The metric function in Eq. (14) satisfies $f(0)=0$ and $f^{\prime}(0)=0$. When analytically extended to negative values of the coordinate $r$, the point $r=0$ acts as an extremal inner horizon, characterized by zero surface gravity. Importantly, the vanishing of the surface gravity prevents classical mass inflation instabilities that are known to plague most regular black hole models [72, 73, 74]. Therefore, an infinite tower of corrections not only cures the singularity, but also has the potential to prevent classical mass inflation instabilities. This situation is similar to that observed in some regular black holes in $2+1$ dimensions [75, 76].
It is important to note that the infinite tower of corrections is essential: truncating the series in Eq. (13) at any finite order would introduce a singularity in the solutions. This can be observed starting from Eq. (13), truncated at $n=n_{\rm max}$, from which we get that near $r=0$ the metric function $f(r)$ behaves as
@@
Regular solutions with $f=-r^{2}/\ell^{2}+\mathcal{O}(r^{3})$ can only be achieved in the limit $n_{\rm max}\to\infty$, provided that $0<\lim_{n\to\infty}(c_{n})^{1/n}<\infty$. This condition ensures that the series has a finite radius of convergence, which we have set to unity without loss of generality.
Discussion. In this work, we have explored one of the most fundamental open problems in theoretical physics – the singularity problem – with far-reaching implications, from the cores of black holes to the very early universe. By including an infinite tower of Lovelock-like corrections to GR through well-defined scalar-tensor Horndeski theories, we have derived solutions that replace the initial singularity with an inflationary phase in the early universe, followed by a graceful exit into standard GR evolution, and solutions that describe regular planar black holes free of singularities. These theories introduce a single additional length scale, $\ell$, which sets the scale for new physics, and crucially, they do not suffer from Ostrogradsky instabilities. Furthermore, the re-summed theories constitute valid Horndeski theories on their own, irrespective of their connection to Lovelock gravity. As a result, the couplings $\{c_{n}\}$ cannot be regarded as fine-tuned in this sense.
The results of this work qualitatively align with those of Ref. [11], which studied higher-dimensional quasi-topological gravity and demonstrated that an infinite series of corrections to GR can act as a mechanism for resolving singularities. Despite being explored in different settings, both Ref. [11] and this work support the paradigm that infinite towers of corrections may be key to the absence of singularities. Such corrections are a common feature of leading candidates for a theory of quantum gravity, including string theory and asymptotically safe quantum gravity. However, our findings suggest that knowledge of the full theory is essential, as truncating the series of corrections typically results in singular spacetimes.
There are several promising directions for future research. A crucial next step would be to generalize our results to black holes with spherical horizons, as these are the most astrophysically relevant solutions. While we have not yet found a way to integrate the field equations in this scenario, an extension of the theories considered here might make this possible – see e.g. the class of theories in Refs. [44, 42], where a slightly different regularization procedure of the quadratic Lovelock invariant is considered. Exploring gravitational collapse scenarios, such as those of Refs. [12, 13], would be an important avenue for further investigation. Additionally, investigating the inflationary predictions of these theories, and their cosmological stability [77] would be an interesting direction. It would also be interesting to study the regular black holes presented in this work from the point of view of thermodynamics and holography [78, 79, 80, 81, 82].

### v2

@@ Abstract
Recent work has shown that regular black holes solutions exist [11] and can form dynamically [12, 13] in higher dimensions, $D\geq 5$, within the framework of quasi-topological gravity [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] when accounting for an infinite tower of curvature corrections. This represents a major step forward in tackling one of the key challenges in theoretical physics: identifying a mechanism to resolve spacetime singularities [25]. Unlike many phenomenological regular black hole models, this approach avoids the need for ad hoc exotic matter or theories that impose fine-tuned relationships between a solution’s integration constants and the theory’s parameters, as often seen in non-linear electrodynamics models [26, 27, 28]. In cosmological settings, it has been shown that, in a similar framework, infinite towers of curvature corrections [23] can replace the initial singularity with an inflationary period [29]. This framework is known as “geometric inflation” [29, 30, 31, 32, 33, 34], see also e.g. Refs. [35, 36, 37] for other frameworks where inflation is achieved purely from gravitational effects. Unfortunately, these theories generally lead to field equations of higher-order in derivatives, except in specific backgrounds, making them susceptible to Ostrogradsky instabilities [38, 39], and quasi-topological gravities do not exist in four-dimensions [18].
@@
In recent years, there has been a surge of interest in the quadratic Lovelock invariant – the Gauss-Bonnet term – and the possibility of a non-trivial Gauss-Bonnet–corrected theory of gravity in four-dimensions. This idea was first introduced in Ref. [42] through a singular rescaling of the Gauss-Bonnet coupling constant in a four-dimensional limit of the higher-dimensional theory. However, this singular limit was later shown to be ill-defined [43], leading to the development of well-defined regularizations. These regularizations resulted in specific four-dimensional Horndeski scalar-tensor theories [44, 45, 46, 47, 48], which preserve many solutions and properties of the higher-dimensional Einstein-Gauss-Bonnet theory. As a result, they have been extensively studied; see Ref. [49] for a comprehensive review.
@@
To obtain Lovelock-like corrections at all orders curvature in four-dimensions, we apply a well-defined dimensional regularization to each Lovelock invariant, resulting in scalar-tensor theories within the Horndeski class [50, 51]. This regularization method was first introduced in Ref. [52] to recover GR-like dynamics in two dimensions by regularizing the Ricci scalar and was later extended in Refs. [44, 45] to formulate a well-defined Gauss-Bonnet theory in four-dimensions. This method relies on the use of two conformally related metrics, $\tilde{g}_{\mu\nu}=e^{-2\phi}g_{\mu\nu}$, and the following limit at each Lovelock order
$\sqrt{-g}\mathcal{L}^{(n)}=\lim_{d\to 2n}\frac{\sqrt{-g}\mathcal{R}^{(n)}-\sqrt{-\tilde{g}}\tilde{\mathcal{R}}^{(n)}}{d-2n},$ (2)
where the expressions inside the limit are to be evaluated in $d$-dimensions, and the limit is taken as $d$ approaches the critical dimension $2n$, where the $n^{\rm th}$ Lovelock invariant becomes topological. This computation for $n=1$ in $2D$, and $n=2$ in $4D$ was performed in detail in Ref. [44], and it can be shown that the limit is well-defined for every $n$, resulting in a scalar-tensor theory, with scalar $\phi$, and second-order equations of motion [53]. Although the limit is taken as $d\to 2n$, the resulting scalar-tensor Lagrangian can be evaluated within a four-dimensional action for any $n$. This four-dimensional theory is inspired by the three-dimensional limits of Gauss-Bonnet gravity and their solutions [54, 55], which, despite not being considered in the Gauss-Bonnet critical dimension $D=4$, lead to solutions similar to those in higher-dimensional Gauss-Bonnet gravity. The final result at each $n$ for the limit (2) is that $\mathcal{L}^{(n)}$ is proportional to a four-dimensional Horndeski Lagrangian
$\mathcal{L}_{H}=G_{2}(\phi,X)-G_{3}(\phi,X)\Box\phi+G_{4}(\phi,X)R$ (3)
$+G_{4X}\left[(\Box\phi)^{2}-\left(\nabla_{\mu}\nabla_{\nu}\phi\right)^{2}\right]+G_{5}(\phi,X)G^{\mu\nu}\nabla_{\mu}\nabla_{\nu}\phi$
$-\frac{G_{5X}}{6}\left[(\Box\phi)^{3}-3\Box\phi\left(\nabla_{\mu}\nabla_{\nu}\phi\right)^{2}+2\left(\nabla_{\mu}\nabla_{\nu}\phi\right)^{3}\right],$
with functions given by [53]
@@
$G_{3}^{(n)}=-2^{n}n(2n-3)X^{n-1},\quad G_{4}^{(n)}=2^{n-1}nX^{n-1},$
[…]
@@
where $\{c_{n}\}$ are dimensionless coupling constants, $\ell$ is a new length scale, and now $\mathcal{L}^{(n)}$ is the Horndeski Lagrangian with functions given by Eq. (4). This theory falls within the shift-symmetric class of Horndeski theories, see e.g. [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66], implying the existence of a conserved current $J^{\mu}$, whose expression is provided in Ref. [65]. The vanishing divergence of this current, $\nabla_{\mu}J^{\mu}=0$, is equivalent to the scalar field equation of motion. Importantly, a sufficient condition to solve the scalar field equation of motion, is $J^{\mu}=0$. The field equations for a generic Horndeski Lagrangian (3) are presented in Refs. [67, 68].
When the whole tower of corrections is considered, the theory described by Eq. (5) can in some cases be re-summed into Horndeski theories with a non-local structure. Examples are presented in the Supplemental Material.
Figure 1: (Left) Time evolution of $a$ and $H^{2}$ (in units $\ell=1$) for the modified Friedmann equation in Eq. (10), and in GR, assuming a radiation dominated universe with initial value for the energy density $\rho=3/(32\pi)$ and scale factor $a=1$ at time $t=1$. In GR, a Big Bang singularity occurs at $t=0$, whereas when an infinite tower of corrections is considered, the singularity is replaced by an inflationary phase. (Right) Slow-roll parameter $\dot{H}/H^{2}$ as a function of the number of e-folds NN (defined such that $N=0$ at the end of inflation, i.e., evolution proceeds from right to left), for the theories with couplings given by $c_{n}=(1-(-1)^{n})/(2n)$, and $c_{n}=1/n!$. The curves are shown up to $N=50$, but they continue all the way to $N\to\infty$.
@@
For the theory (5) on an FLRW background, the only non-trivial component of the shift-symmetry current is $J^{t}$. The scalar field equation can be integrated to the form $J^{t}=Q/a^{3}$, where $Q$ is an integration constant interpreted as the shift-symmetry scalar charge, and $J^{t}$ is interpreted as the scalar charge density. At each order $n$, the current vanishes if the scalar field obeys
@@
where $a_{0}$ is an arbitrary integration constant. This profile solves the scalar field equation of motion for any $n$, implying $Q=0$. While we were unable to find exact solutions for $Q\neq 0$, solutions with a vanishing charge/current are well-motivated for several reasons: (i) they are attractors in an expanding universe; (ii) the scalar adopts the profile (7), which aligns with expectations from the higher-dimensional origin of the theory, where it represents the logarithm of the scale factor of the extra-dimensional space prior to taking the 4D limit [46, 47]; (iii) the absence of additional charges is expected when interpreting the theory as a 4D Lovelock theory; and (iv) as shown below, it leads to Friedmann equations that match those of Lovelock gravity [69].
Using the scalar field profile (7) in the Einstein equations, we obtain the generalized Friedmann equation
@@
As in Lovelock gravity [69], the $n^{\text{th}}$ order contribution to the Friedmann equation is proportional to $H^{2n}$. As an example, for the theory with $c_{n}=\left(1-(-1)^{n}\right)/(2n)$, we obtain $F(H^{2})=\tanh^{-1}\left(\ell^{2}H^{2}\right)/\ell^{2}$ such that the Friedmann equation can be written as
@@
Examples for other choices of couplings are presented in the Supplemental Material. In this scenario, as $\rho\to\infty$ in the very early universe [^2] From the continuity equation, we obtain $\rho=\rho_{0}a^{-3(1+\omega)}$, where $\omega$ is the equation of state parameter, $p=\omega\rho$. In the early universe, radiation ($\omega=1/3$) dominates over matter ($\omega=0$)., the spacetime must asymptotically approach a non-singular de Sitter geometry with $H^{2}\to 1/\ell^{2}$. Consequently, the Big Bang singularity is replaced by an inflationary phase in the early universe. This non-singular scenario differs from others, such as bounces, as there is no contracting phase, and the energy density is not bounded from above. However, infinite energy density is never attained in finite time. This inflationary phase can be observed in Fig. 1 (left) we compare the time evolution of $H^{2}$ and $a$ for a radiation dominated universe [^3] The same qualitative conclusions hold when other initial conditions, and equations of state, such as matter ($\omega=0$), are considered. in GR, and in the theory with Friedmann equation given in Eq. (10). For late times, the evolution of the universe is well-described by GR. Other couplings we have explored, such as $c_{n}=1$, $c_{n}=1/n$, $c_{n}=1/n!$, and $c_{n}=1/(n-1)!$, also replace the initial singularity with an inflationary phase. In Fig. 1 (right), we plot the slow-roll parameter $\dot{H}/H^{2}$ as a function of the number of e-folds $N$, for the theories with $F(H^{2})=\tanh^{-1}\left(\ell^{2}H^{2}\right)/\ell^{2}$ and $F(H^{2})=(\exp\left(\ell^{2}H^{2}\right)-1)/\ell^{2}$. We observe an inflationary period ($\dot{H}/H^{2}<1$) followed by a graceful transition into a radiation dominated universe ($\dot{H}/H^{2}=2$). As shown in Refs. [29, 73], the singularity is avoided provided the infinite tower is included, with the scale factor exhibiting at least exponential growth in the far past. Thus, both singularity resolution and inflation emerge as generic features, largely insensitive to the specific couplings $c_{n}$ and the choice of initial conditions.
Regular black holes. Black holes are predicted to form through gravitational collapse. The simplest case, initially explored by Oppenheimer and Snyder [74], involves the collapse of a homogeneous and isotropic sphere of pressureless dust. The interior metric describing this collapsing dust, which must be matched to an exterior metric, can be represented by the FLRW line-element and its dynamics governed by the Friedmann equations. In GR, the dust collapses in finite time to form a singularity and a Schwarzschild black hole. However, when the full tower of corrections in Eq. (5) is taken into account, the Friedmann equations are modified. For the examples studied in the previous section, such as Eq. (10), the collapsing dust, in the marginally-bound case [75], never forms a singularity in finite time. Beyond the marginally-bound case, the interior geometry is described by a spatially curved FLRW metric. In the Supplemental Material, we present the modified Friedmann equations of the theory (5) in this setting. Even in this case, $H$ remains bounded for the cases we studied, and the conclusions drawn from the spatially-flat case apply. This raises questions about whether singularities can form at all through gravitational collapse in this framework.
@@
where the function $F(x)$ was defined in Eq. (9) and $f_{\rm GR}(r)$ in Eq. (12). This leads us to the conclusion that the class of theories examined in this letter satisfy a staticity theorem [^5] To our knowledge, this is the first example of a staticity theorem holding for a class of higher-derivative theories beyond GR in four-dimensions. However, see Refs. [92, 104, 105, 13, 12] for specific cases in $2+1$ dimensions and higher-dimensional scenarios where staticity (and in some cases, Birkhoff) theorems have been established.: assuming the scalar field profile (13), any solution of the form (11) within the theory (5) must be both static and unique, characterized by $N(t,r)=1$ and determined by the solution to Eq. (14). This result can be used to study the collapse of matter and the dynamical formation of black holes, following e.g. the approach of Refs. [13, 12]. Note that the vanishing of $J^{\mu}$ is necessary, as the scalar field equation takes the form $J^{r}=Q/r^{2}$, and if $Q\neq 0$, the scalar quantity $J^{\mu}J_{\mu}=Q^{2}/(r^{4}f(r))$ diverges at the horizon, and at $r=0$.
As discussed in Refs. [82, 83, 84, 85], for static spacetimes of the form (11), a necessary and sufficient condition for the regularity of all independent components of the Riemann tensor, $R^{\mu\nu}_{\phantom{\mu}\phantom{\nu}\alpha\beta}$, and consequently all curvature invariants constructed from it, is that the Kretschmann scalar remains finite everywhere. By analyzing the Kretschmann scalar, one can show that the singularity at $r=0$ is resolved if the metric function behaves as $f(r)=-r^{2}/\lambda^{2}+\mathcal{O}(r^{3})$ near the origin, where $\lambda$ is a constant with dimensions of length. From Eq. (14), we find that resolving the singularity is possible if the function $F(x)$ develops a pole as $x\to 1/\lambda^{2}$. Sufficient conditions for the existence of a pole are: $c_{n}\geq 0$ for all $n$; a non-zero finite radius of convergence $\mathfrak{R}=1/\lim_{n\to\infty}(c_{n})^{1/n}$; the series $\sum_{n}c_{n}\mathfrak{R}^{n}$ diverges. This class of couplings resolve the singularity of both cosmological models and black holes. Since there are infinitely many possible choices of the coefficients $\{c_{n}\}$ that could induce such a pole in $F(x)$, we adopt as a representative example the same couplings used in the previous section, $c_{n}=\left(1-(-1)^{n}\right)/(2n)$, which leads to a pole at $x\to 1/\ell^{2}$. In this case, the solution to Eq. (14) is
@@
where the GR solution is recovered in the limit $\ell\to 0$. Additional examples are provided in the Supplemental Material. The location of the event horizon remains unchanged from the GR case, $r_{H}=(2L^{2}M)^{1/3}$, and the solution is asymptotically AdS. This metric function is everywhere smooth and non-divergent, even when analytically extended to negative values of the coordinate $r$ [86, 87]. The solution is also free of standard curvature singularities as the Kretschmann scalar is everywhere bounded, obeying $\lim_{r\to 0}R_{\mu\alpha\nu\beta}R^{\mu\alpha\nu\beta}=24/\ell^{4}$. Additional solutions which are non-singular at $r=0$ can be found, for instance, in the theories where $c_{n}=1$ or $c_{n}=1/n$.
The metric function in Eq. (15) satisfies $f(0)=0$ and $f^{\prime}(0)=0$. Thus, $r=0$ is a Killing horizon associated to the Killing vector $\xi^{\mu}=(\partial_{t})^{\mu}$, since $g_{\mu\nu}\xi^{\mu}\xi^{\nu}|_{r=0}=-f(0)=0$. Moreover this horizon is extremal in the sense that it has zero surface gravity, $\kappa|_{r=0}=-f^{\prime}(0)/2=0$. Details on the causal structure of the spacetime (15) are given in the Supplemental Material. The vanishing of the surface gravity prevents classical mass inflation instabilities that plague most regular black hole models [88, 89, 90]. Thus, an infinite tower of corrections not only cures the singularity, but also has the potential to prevent classical mass inflation instabilities. This situation is similar to that observed in some regular black holes in $2+1$ dimensions [91, 92].
The infinite tower of corrections is essential: truncating the series in Eq. (14) at any finite order introduces a singularity in the solutions. This can be observed starting from Eq. (14), truncated at $n=n_{\rm max}$, from which we get that near $r=0$ the metric function $f(r)$ behaves as
@@
Regular solutions with $f=-r^{2}/\ell^{2}+\mathcal{O}(r^{3})$, can only be achieved in the limit $n_{\rm max}\to\infty$, provided that $\mathfrak{R}$ is finite.
Discussion. In this work, we have explored one of the most fundamental open problems in theoretical physics – the singularity problem. By including an infinite tower of Lovelock-like corrections to GR through well-defined scalar-tensor Horndeski theories, we have derived solutions that replace the initial singularity with an inflationary phase in the early universe, followed by a graceful exit into standard GR evolution, and solutions that describe regular planar black holes free of singularities. These theories introduce a single additional length scale, $\ell$, which sets the scale for new physics, and crucially, they do not suffer from Ostrogradsky instabilities.
The results of this work qualitatively align with those of Ref. [11], which studied higher-dimensional quasi-topological gravity and showed that an infinite series of corrections to GR can act as a mechanism for resolving singularities. Despite being explored in different settings, both Ref. [11] and this work support the paradigm that infinite towers of corrections may be key to the absence of singularities. Such corrections are a common feature of leading candidates for a theory of quantum gravity. However, our findings suggest that knowledge of the full theory is essential, as truncating the series of corrections typically results in singular spacetimes.
There are several promising directions for future research. A crucial next step would be to generalize our results to black holes with spherical horizons, as these are the most astrophysically relevant solutions. While we have not yet found a way to integrate the field equations in this case, an extension of the theories considered here might make this possible – see e.g. the class of theories in Refs. [48, 46], where a slightly different regularization procedure of the quadratic Lovelock invariant is considered. Exploring gravitational collapse scenarios, such as those of Refs. [12, 13] – see also Refs. [93, 94, 95] – would be an important avenue for further investigation. Additionally, investigating the inflationary predictions of these theories, and their cosmological stability – with respect to both spatial anisotropies [96], and to perturbations [97, 98] – would be an interesting direction. The generic no-go theorems of Horndeski gravity [97, 98] are expected to apply in our case. Nevertheless, our construction can be extended beyond this framework to circumvent these theorems. For instance, introducing a distinct scalar field at each order in the Lovelock expansion yields the same phenomenology for both cosmology and black holes as the single scalar case studied in this work. It would also be interesting to study the black holes presented in this work from the point of view of thermodynamics and holography [99, 100, 101, 102, 103].